‘Bhartdarshan’ is an Internet-based travel agency wherein customers can see videos of the


cities they plan to visit. The number of hits daily is a normally distributed random variable


with a mean of 10,000 and a standard deviation of 2,400.


a. What is the probability of getting more than 12,000 hits?


b. What is the probability of getting fewer than 9,000 hits?

1
Expert's answer
2022-05-12T09:19:57-0400

We have a normal distribution, "\mu=10000, \sigma=2400."

Let's convert it to the standard normal distribution,

"z=\cfrac{x-\mu}{\sigma}.\\
\text{a. } z_1=\cfrac{12000-10000}{2400}=0.83;\\
P(X>12000)=P(Z>0.83)=\\
=1-P(Z<0.83)=\\
=1-0.7967=0.2033
\text{ (from z-table)}.\\
\text{b. } z_2=\cfrac{9000-10000}{2400}=-0.42;\\
P(X<9000)=P(Z<-0.42)=\\
=0.3372
\text{ (from z-table)}."